# From reductionism to realism: Holistic mathematical modelling for complex biological systems

### v1

@@ III-C Automatic modelling
Whilst we have advocated for a holistic MM that embraces complexity, advances in simulation-based science open the door for automatic modelling (AM). Data-driven approaches to MM typically extract a low number of important latent features from high-dimensional, complex systems. In an approach akin to model selection, AM will be able to leverage simulations and ML techniques to perform a posteriori reductionism. Within this framework, modellers initially construct detailed models that incorporate all available complexities. Subsequently, employing optimisation strategies from computational statistics and ML, these models are systematically simplified by fitting them to experimental data while simultaneously penalising model complexity. This approach allows AM to extract, in a sparse manner, the most relevant features or underlying physical principles essential for accurately explaining the observed system behaviour.
This framework remains to be formally established, but its alignment with existing model-selection and regularisation techniques suggests that it will become feasible within the coming decades. One related approach is that of the Bayesian machine scientist [72], which explores and selects between candidate models using Monte-Carlo methods and Bayesian inference, penalising model-complexity, in order to perform symbolic regression. AM differs significantly from current model-reduction and coarse-graining methods [73] as well as traditional system identification and data-driven discovery approaches. Rather than relying on subjective decisions by modellers or opaque nonlinear transformations, this method objectively identifies essential features. Consequently, it transparently reveals, in an interpretable and physically meaningful manner, the dominant features responsible for explaining experimental data through an automated, mechanistic process.
@@ IV Modelling as an inverse-problem
where $\mathbf{f}$ represents the mechanistic dynamics of the system, known as the drift, and $\xi$ represents the additive noise, known as the diffusion. The goal of a data-driven method is to infer $\mathbf{f}$ from the MVTS [^4] Stochastic methods often also attempt to infer the diffusion tensor $\mathbf{D}(\mathbf{x},t)=\frac{1}{2}\xi(\mathbf{x},t)\xi^{\top}(\mathbf{x},t)$ [76]. [76]. Here methods can vary, with some attempting to infer the exact mathematical terms in $\mathbf{f}$, typically known as symbolic regression methods, and others approximating $\mathbf{f}$ with deep learning methods. Popular exisiting techniques include SINDy [77], which performs symbolic regression for nonlinear, deterministic dynamics, as well as its extensions to stochastic differential equations (SDEs) [78] and PDE [79] models. AI-Feynman is one deep-learning approach for discovering physical laws directly from observed data [80]. For SDEs, a number of other methods exist such as stochastic force inference [81, 82], InferenceMAP [83] and maximum-likelihood estimation [76]. These methods have particular applications to Brownian dynamics in soft-matter systems [84, 85, 86]. More recently, extensions have been considered for high-dimensional networked systems [87, 88].
Beyond symbolic inference methods, the dynamics, $\mathbf{f}$, can be approximated using generative machine learning [89]. Deep learning architectures such as recurrent neural networks [^5] Reservoir computers [90] are also often used for predicting nonlinear dynamics, however they are a form of RNN. (RNNs) and transformers have been used to learn and predict dynamics from observations with applications to neural and molecular data [91, 92, 93].
@@ IV-A Latent representations for dynamical systems
Opposite to dimensionality reduction, ‘lifting’, projection of dynamics into higher-dimensional space, is also particular useful for solving inverse problems. In particular, the Koopman operator lifts nonlinear dynamics into an infinite-dimensional system where the dynamics are linear [107]. Finite approximations of the Koopman operator can be learned using dynamic mode decomposition [108] and deep learning [109], as illustrated in Panel B, Fig. 5.
@@ V A fundamental physics for biology
We have argued that theoretical biology should turn towards complex, empirical models with predictive capabilities, even if these models occasionally complicate traditional mathematical analyses. Nevertheless, this shift does not negate the importance of identifying fundamental theoretical principles that govern biological systems. Unlike physics, biology still lacks clear translational, foundational principles. Thus, the field urgently requires the development of a ’theoretical physics for living systems’. Whilst significant progress remains limited, several promising directions have emerged in recent years.
@@
Models of collective phenomena often exhibit phase-transitions occurring at critical points [110]. In 1987, Bak et al. hypothesised that complex systems may self-organise to the critical point, where both correlation-length across the system and sensitivity to perturbations are maximised [111, 112]. This theory sparked particular interest in biological sciences, where evidence of self-organised criticality (SOC) has been found in neuronal firing [113] and genetic expression [114, 115], leading to more general theories about how critical dynamics facilitate biological function [116, 117]. Moreover, recent work formalising the concept of Griffith’s phases, parameter regions showing critical-like dynamics as opposed to isolated points, promise new avenues for the furtherance of this theory [118], which could prove to be both a governing principle for biological systems, and explain observed biological complexity.
@@
Another prominent theory accounting for self-organisation in biological systems is the free-energy principle (FEP) [119]. Originally proposed as a theory of computation in the brain, the FEP suggests that biological organisms encode sensory information in the form of a probabilistic world model that is updated according to Bayesian mechanics [^6] It is worth noting that the ‘energy’ in the FEP is not usually a physical energy but a larger unifying quantity that plays a similar role. Next, it posits that organisms act in order to minimise the surprise [^7] The negative log probability of an outcome. i.e. the difference between sensory information and the world model. Whilst the FEP is formulated to account for perception-actions loops in the brain, it represents a more general theory of decision making that has been recently used to model collective behaviour in swarms [120]. More generally, this theory, amongst other theories of agentic decision making, can be used to construct complex interacting models that begin from a cognitive, mechanistic basis.
@@
Nonequilibrium thermodynamics has long been regarded as central to an understanding of the energetic constraints of biological processes [14, 121]. However, the theoretical formalism for describing the stochastic thermodynamics of fluctuating, microscopic processes is particularly recent [122, 123]. Subsequently, stochastic thermodynamics has become extremely useful for describing the activity of biological processes, inspiring a plethora of techniques for analysing the thermodynamics of observed biological activity [81, 124, 125, 126, 127], with recent interest in neural activity [128, 129, 130]. Despite these advancements, there remain a number of open questions regarding the thermodynamic limits of complexity and information processing [131], and how this pertains to biological processes. Nevertheless, investigating the thermodynamic constraints on biological processes promises to be an avenue ripe for discovery.
@@
Finally, in an effort to describe the mechanisms by which complex structures, such as amino acids, evolve, Sharma et al. recently proposed assembly theory (AT) [132]. AT quantifies the complexity of a structure in terms of the amount of selection required to produce it. Whilst theoretical in nature, AT offers an avenue for understanding the constraints on the complexity of molecules, giving insight into evolution, but also yielding a testable indicator for the existence of extraterrestrial life [133].
@@
In the near future, mathematical biology and complex systems should prioritise developing empirically predictive and practical models, while continuing the pursuit of underlying theoretical principles. This area of investigation requires bold theorising that, whilst controversial at times, promises to move the field forward. Modellers should err towards theoretical predictions that are easily testable on existing data, and are rooted in established and well-tested physical principles. Progress on this front will only serve to improve the mechanistic understanding and empirical usefulness of the complex models that we advocate for. Again, these goals may be symbiotic. As mathematical and computational models begin to yield useful, actionable insights, the apparent resistance in empirical biological sciences to theoretical hypotheses may shift, ultimately blurring the edges between domains and leading to a truly interdisciplinary approach.
@@ VI Back to the future
Nonetheless, we reaffirm that the identification of fundamental theoretical principles should remain the ultimate goal of biology. We highlight recent key advances offering translational, foundational insights into longstanding biological questions, building upon established theories from physics.

### v2

@@ III-C Automatic modelling
Advances in simulation-based science open the door for a new paradigm that we call automatic modelling (AM). By combining techniques in ML, system identification, and model selection, AM is an encompassing term for methods that are able to leverage simulations to perform ‘theory-discovery’ alongside a posteriori reductionism. We envisage a framework where modellers initially construct detailed models that incorporate all available complexities. Subsequently, employing optimisation strategies from computational statistics and ML, these models are systematically simplified by fitting them to experimental data while simultaneously penalising model complexity. This approach would allow AM to extract, in a sparse manner, the most relevant features or underlying physical principles essential for accurately explaining the observed system behaviour. Depending on the required predictive power, users of AM could adjust model complexity against model interpretability. It would answer the crucial question: how much of the data can be explained by a reduced set of key mechanisms?
Whilst this framework remains to be formally established, it shows strong alignment with existing model-selection and regularisation techniques, thus suggesting that it will become feasible within the coming decades, especially as LLM prompts for scientific discovery become increasingly conversational. Moreover, there exist some preliminary related approaches such as that of the Bayesian machine scientist [81], which explores and selects between candidate models using Monte-Carlo methods and Bayesian inference, penalising model-complexity, in order to perform system identification. AI-Feynman [82], AI-Descartes [83], and AI-Hilbert [84] are three related approaches that attempt to discover physical laws directly from observed data by blending theory and experimental data using ML techniques. AM represents a future ideal where rather than relying on subjective decisions by modellers or opaque nonlinear transformations, the underlying engine could objectively identify models with essential features. This would reveal the dominant mechanism responsible for explaining experimental data through an automated process.
III-D Generative ML and AI
It would be remiss to discuss the future of MM without making reference to the incredible developments in generative forms of ML and AI. In particular, innovations such as the transformer architecture [85], derivatives such as large language models (LLM), or other techniques, such as diffusion models [86], seem posed to revolutionise society, including science, even after taking into account the unreasonable media hype. In particular, generative ML stands in stark contrast with the discriminative models of the past, as these methods learn a dynamical process that is capable of sampling new candidates, including protein structures [87], molecules [88], and human-level language capabilities [89].
Despite their generative and predictive power, these uninterpretable, black-box models seem at odds with MM, nevertheless there are a number of complementary directions. Whilst transformers lack the biological plausibility of spiking or recurrent networks, their function resembles that of grid-cells in the hippocampus [90], suggesting that they could yet teach us something about the brain [91], or simply prompt new methods in dimensionality reduction and visualisation for biological data [92]. Moreover, transformers have lead to the development of so called foundation models, which can subsequently be fine-tuned for a number of scientific tasks, as diverse as predicting chaotic [93] and stochastic dynamics [94] or analysing single-cell data [95]. As such algorithms develop, they become a new tool in the arsenal of the modern mathematical modeller, opening up a plethora of synergistic methods [96], including improved model calibration or data-driven approaches to dynamical systems.
@@ IV Modelling as an inverse-problem
where $\mathbf{f}$ represents the mechanistic dynamics of the system, known as the drift, and $\xi$ represents the additive noise, known as the diffusion. The goal of a data-driven method is to infer $\mathbf{f}$ from the MVTS [^6] Stochastic methods often also attempt to infer the diffusion tensor $\mathbf{D}(\mathbf{x},t)=\frac{1}{2}\xi(\mathbf{x},t)\xi^{\top}(\mathbf{x},t)$ [99]. [99]. Here methods can vary, with some attempting to infer the exact mathematical terms in $\mathbf{f}$, typically known as sparse regression methods, and others approximating $\mathbf{f}$ with deep learning methods. Popular exisiting techniques include SINDy [100], which performs sparse regression for nonlinear, deterministic dynamics, as well as its extensions to stochastic differential equations (SDEs) [101] and PDE [102] models. For SDEs, a number of other methods exist such as stochastic force inference [103, 104], InferenceMAP [105] and maximum-likelihood estimation [99]. These methods have particular applications to Brownian dynamics in soft-matter systems [106, 107, 108, 109]. More recently, extensions have been considered for high-dimensional networked systems [110, 111].
Beyond sparse force inference methods, the dynamics, $\mathbf{f}$, can be approximated using generative machine learning [112]. Deep learning architectures such as recurrent neural networks [^7] Reservoir computers [113] are also often used for predicting nonlinear dynamics, however they are a form of RNN. (RNNs) and transformers have been used to learn and predict dynamics from observations with applications to neural and molecular data [114, 94, 115].
@@ IV-A Latent representations for dynamical systems
Opposite to dimensionality reduction, ‘lifting’, projection of dynamics into higher-dimensional space, is also particular useful for solving inverse problems. In particular, the Koopman operator lifts nonlinear dynamics into an infinite-dimensional system where the dynamics are linear [129, 130, 131]. Finite approximations of the Koopman operator can be learned using dynamic mode decomposition [132] and deep learning [133], as illustrated in Panel B, Fig. 5.
@@ V A fundamental physics for biology
We have argued that theoretical biology should turn towards complex, empirical models with predictive capabilities, even if these models occasionally complicate traditional mathematical analyses. Nevertheless, this shift does not negate the importance of identifying fundamental theoretical principles that govern biological systems, but instead, supports it. Unlike physics, biology still lacks clear translational, foundational principles. Thus, the field urgently requires the development of a ’theoretical physics for living systems’ [134]. Whilst significant progress remains limited, several promising directions have emerged in recent years, where holistic MM will play a crucial role.
@@
Models of collective phenomena often exhibit phase-transitions occurring at critical points [135]. In 1987, Bak et al. hypothesised that complex systems may self-organise to the critical point, where both correlation-length across the system and sensitivity to perturbations are maximised [136, 137]. This theory sparked particular interest in biological sciences, where evidence of self-organised criticality has been found in neuronal firing [138] and genetic expression [139, 140], leading to more general theories about how critical dynamics facilitate biological function [141, 142]. Moreover, recent work formalising the concept of Griffith’s phases, parameter regions showing critical-like dynamics as opposed to isolated points, promise new avenues for the furtherance of this theory [143], which could prove to be both a governing principle for biological systems, and explain observed biological complexity. The study of criticality in neural systems has motivated a number of data-driven approaches to dynamical systems, in particular to the problem of inferring collective properties from a fraction of total observables [144, 145] - techniques with implications far beyond neuroscience.
@@
Another prominent theory accounting for self-organisation in biological systems is the free-energy principle (FEP) [146]. Originally proposed as a theory of computation in the brain, the FEP suggests that biological organisms encode sensory information in the form of a probabilistic world model that is updated according to Bayesian mechanics [^8] It is worth noting that the ‘energy’ in the FEP is not usually a physical energy but a larger unifying quantity that plays a similar role.. Next, it posits that organisms act in order to minimise the surprise [^9] The negative log probability of an outcome. i.e. the difference between sensory information and the world model. Whilst the FEP is formulated to account for perception-actions loops in the brain, it represents a more general theory of decision making that has been recently used to model collective behaviour in swarms [147]. More generally, this theory, amongst other theories of agentic decision making, can be used to construct complex interacting models, such as ABMs and DT, that begin from a cognitive, mechanistic basis.
@@
Nonequilibrium thermodynamics has long been regarded as central to an understanding of the energetic constraints of biological processes [15, 148]. However, the theoretical formalism for describing the stochastic thermodynamics of fluctuating, microscopic processes is particularly recent [149, 150]. Subsequently, stochastic thermodynamics has become extremely useful for describing the activity of biological processes, inspiring a plethora of techniques for analysing the thermodynamics of observed biological activity [151, 152, 153, 154], with recent interest in neural activity [155, 156, 157]. In particular, a major focus has been placed on the reconstruction of force fields from stochastic trajectories, to estimate thermodynamic quantities [103, 158]. In addition, recent work has shown that ML algorithms can learn thermodynamic laws, such as the arrow of time, opening new avenues for the analysis of biological data [159]. Despite these advancements, there remain a number of open questions regarding the thermodynamic limits of complexity and information processing [160], and how this pertains to biological processes. Nevertheless, investigating the thermodynamic constraints on biological processes promises to be an avenue ripe for discovery.
@@
Finally, in an effort to describe the mechanisms by which complex structures, such as amino acids, evolve, Sharma et al. recently proposed assembly theory [161]. Assembly theory quantifies the complexity of a structure in terms of the amount of selection required to produce it. Whilst theoretical in nature, assembly theory offers an avenue for understanding the constraints on the complexity of molecules, giving insight into evolution, but also yielding a testable indicator for the existence of extraterrestrial life [162]. Moreover, its algorithmic nature makes it well suited for simulation-based discovery and in-silico experimentation, ultimately offering new directions towards artificial or digital life [163].
@@
In the near future, mathematical biology and complex systems should prioritise developing empirically predictive and practical models, while continuing the pursuit of underlying theoretical principles. This area of investigation requires bold theorising that, whilst controversial at times, promises to move the field forward. Modellers should err towards theoretical predictions that are easily testable on existing data, and are rooted in established and well-tested physical principles. Progress on this front will only serve to improve the mechanistic understanding and empirical usefulness of the complex models that we advocate for. Again, these goals may be symbiotic. As mathematical and computational models begin to yield useful, actionable insights, the apparent resistance in empirical biological sciences to theoretical hypotheses may shift, ultimately blurring the edges between domains and leading to a truly interdisciplinary approach. Moreover, the holistic MM advanced here will naturally lead to the development of further theories. For example, system identification will reveal novel mechanisms governing biological processes, whilst simulation-based approaches amount to in-silico experimentation where theories can be efficiently tested and remoulded. Together, these techniques yield a novel approach to science that is better fit for both our unsolved problems and the modern world.
@@ VI Back to the future
Nonetheless, we reaffirm that the identification of fundamental theoretical principles should remain the ultimate goal of biology. We highlight recent key advances offering translational, foundational insights into longstanding biological questions, building upon established theories from physics. We believe that holistic mathematical modelling promises to yield new paradigms capable of constructing and testing biological theories efficiently, hence achieving the dream of a truly quantitative theory of life.